1. You throw a ball purely horizontally through the air, it bounces off a wall and continues at about the same speed back towards you also purely horizontally (pretending gravity isn't a factor).
Then can we say that the impulse delivered onto the ball from the collision with the wall is about double the momentum it had at the start?

True

False

2. The total momentum of a system subject to only internal forces (one that is isolated) is conserved.

True

False

3. Which of the following is an example of an elastic collision.

A.

A person running and jumping onto a stationary skateboard, coasting away.

B.

A bullet hitting a block of wood and embedding itself into the wood.

C.

A gun recoiling from firing a bullet.

D.

A bowling ball hitting a bowling pin.
Answer #1

suppose initial velocity = v m/s

final speed but in reverse direction = -v m/s

impulse =ΔP = Pf-Pi = m* ( vf-vi ) = m* ( v--v) =2 *mv so statement is correct

True

2)

True

3)

D.

A bowling ball hitting a bowling pin